Determine the lateral area of a cylinder of height 300mm and base radius of 100mm.
1
Expert's answer
2020-12-08T10:13:31-0500

"h = 300\; mm \\

r = 100 \;mm \\

Lateral \;area\; of \;cylinder = 2πrh \\

= 2 \times 3.14 \times 100 \times 300 \\

= 188520 \; mm^2"

Answer: 188520 mm2